### v1

# CoherenDream: Boosting Holistic Text Coherence in 3D Generation via Multimodal Large Language Models Feedback

Abstract
Score Distillation Sampling (SDS) has achieved remarkable success in text-to-3D content generation. However, SDS-based methods struggle to maintain semantic fidelity for user prompts, particularly when involving multiple objects with intricate interactions. While existing approaches often address 3D consistency through multiview diffusion model fine-tuning on 3D datasets, this strategy inadvertently exacerbates text-3D alignment degradation. The limitation stems from SDS’s inherent accumulation of view-independent biases during optimization, which progressively diverges from the ideal text alignment direction. To alleviate this limitation, we propose a novel SDS objective, dubbed as Textual Coherent Score Distillation (TCSD), which integrates alignment feedback from multimodal large language models (MLLMs). Our TCSD leverages cross-modal understanding capabilities of MLLMs to assess and guide the text-3D correspondence during the optimization. We further develop 3DLLaVA-CRITIC - a fine-tuned MLLM specialized for evaluating multiview text alignment in 3D generations. Additionally, we introduce an LLM-layout initialization that significantly accelerates optimization convergence through semantic-aware spatial configuration. Comprehensive evaluations demonstrate that our framework, CoherenDream, establishes state-of-the-art performance in text-aligned 3D generation across multiple benchmarks, including T3Bench and TIFA subset. Qualitative results showcase the superior performance of CoherenDream in preserving textual consistency and semantic interactions. As the first study to incorporate MLLMs into SDS optimization, we also conduct extensive ablation studies to explore optimal MLLM adaptations for 3D generation tasks.
Figure 1: While original Score Distillation suffer from per-view bias accumulation resulting in objects omission, Textual Coherent Score Distillation first incorporates feedback of MLLM into SDS optimization, leading to textual consistent 3D results.
1 Introduction
3D content generation is essential for diverse applications, including gaming, virtual reality, and robotics simulation. Recently, significant progress has been made in text-to-3D generation through Score Distillation Sampling (SDS) [30, 41, 21, 4, 34, 26]. SDS-based methods enable high-quality and diverse 3D generation based on user-provided text inputs, effectively distilling image distributions from 2D diffusion models [32] into parameterized 3D representations.
Despite these advances, significant challenges persist in generating 3D content that faithfully adheres to user prompts, especially those involving multiple objects and intricate interactions. The primary limitation stems from SDS’s per-view distillation mechanism, which lacks explicit constraints for maintaining global textual consistency across multiple updates. Per-view independent optimization accumulates biases during the optimization process, causing a progressive deviation from the intended text-to-3D alignment. Furthermore, while recent variants of SDS [34, 24, 18] adopt fine-tuned diffusion models on specific 3D datasets to enhance 3D consistency, they exacerbate the problem of textual inconsistency in 3D generations [15, 18], leading to object omissions and physically implausible spatial relationships, as illustrated in Fig. 1.
How to develop coherent text-to-3D generation remains relatively unexplored. Recent research efforts, such as JointDreamer [15], emphasize the importance of maintaining textual consistency within the original diffusion model compared to its variants [34]. DreamView [44] focuses on training a view-specific text-to-diffusion model to enhance alignment with specific viewpoints. However, both approaches still face the challenge of bias accumulation during SDS optimization. Another intuitive way to alleviate textual inconsistency in 3D generation is through compositional strategies, such as GALA3D [46] and GraphDreamer [9]. Nevertheless, GALA3D [46] suffers from non-overlap decomposition, which leads to unnatural spatial relationships. Although GraphDreamer [9] advances the modeling of object relationships through scene graphs, it remains limited by compositional optimization, resulting in implausible object combination. In contrast, our work aims to generate coherent results within a holistic 3D representation where object relationships and spatial arrangements maintain textual consistency across all viewpoints.
Recent advances in multimodal large language models (MLLMs) demonstrate powerful cross-modal understanding capabilities, leading to successful integrations in text-to-image generation  [37, 12, 8, 19, 5]. While 2D counterparts are promising, directly adapting MLLMs to 3D generation is not feasible, owing to their limited grasp of 3D representation and spatial relationships. We address these limitations by reformulating text-3D alignment as across-view question-answering tasks. Our key insight is that MLLMs’ semantic reasoning capabilities can complement gradient updates from image distribution when properly contextualized within SDS optimization. To this end, we present CoherenDream, which firstly regards MLLMs as dynamic semantic assessor to ensure textual consistency in SDS optimization. We encode the input text and the across-view image distribution from the diffusion model as text-format ground truth, measuring the loss against the text sequences predicted by the MLLM to serve as feedback. This introduces Textual Coherent Score Distillation (TCSD), which adopts MLLM feedback in the SDS optimization to steer the optimization toward a textual-consistent distribution.
Current MLLMs are primarily designed as language assistants, they lack the proficiency to critique 3D generation effectively. To bridge this gap, we further develop a 3DLLaVA-CRITIC based on LLaVA-OneVision 0.5b [17] to enhance the quality of the feedback. Specifically, we design a view-aware data collection pipeline to simulate the gradient updates during SDS optimization, ensuring the delivery of accurate feedback. Furthermore, we introduce LLM-layout initialization, which is integrated with TCSD to warm up the 3D representation, thereby enhancing the textual consistency for image distribution from diffusion models. Extensive experimental results demonstrate that CoherenDream not only generates text-coherent 3D content but also outperforms other text-to-3D generation methods in terms of quality and quantitative metrics.
In summary, our contributions are as follows:
  • We introduce a novel Textual Coherent Score Distillation (TCSD) for text-coherent 3D generation, guiding optimization with MLLM feedback.
  • We propose a view-aware data collection pipeline and fine-tune the 3DLLaVA-CRITIC to provide accurate feedback between text and 3D representations.
  • Our CoherenDream establishes a new benchmark in coherent text-to-3D generation, producing 3D content that faithfully reflects user inputs.
2 Related Works
SDS-based Text-to-3D Generation.
The Score Distillation Sampling (SDS) algorithm has achieved surprising results in text-to-3D generation. It pioneers by [30], utilizing 2D diffusion model priors [32] to optimize 3D representations. Recent advancements have further refined this technique by enhancing 3D representations [21, 4, 45, 39], improving generation quality [14, 41, 47, 20], and ensuring 3D consistency [15, 34, 18, 33, 2]. Despite these impressive results, these methods still struggle with multi-object prompts [10] and semantic interaction. These challenges often stem from the lack of global textual consistency and accumulation of view-independent biases. To address it, we introduce MLLM feedback into SDS, which dynamically revises the update direction of 3D representations, guiding the score distillation process toward text-3D alignment.
Figure 2: The overview of CoherenDream framework. CoherenDream involves LLM-Layout initialization, semantic coherent score distillation and 3DLLaVA-CRITIC with three kinds of guidance tasks, producing text-3D aligned results from MLLM feedback.
Compositional Text-to-3D Generation.
One intuitive approach for text-aligned 3D generation is decomposing holistic representations into individual components for separate optimization before recombination. However, existing methods [22, 3] suffer from quality degradation due to the challenges in managing layout constraints during NeRF optimization and their dependence on potentially inaccurate predefined layouts. Recently, GALA3D [46] attempts to use 3D Gaussians representation while dynamically refine layout during generation. However, its non-overlapping constraint leads to unnatural spatial relationships and scale inconsistencies (e.g., excessive spacing between the girl and dinosaur in Fig. 4). GraphDreamer [9] advances the field by incorporating automatically parsed scene graphs to guide object interactions. Nevertheless, it inherits the fundamental limitations of compositional approaches, struggling to maintain holistic coherence, particularly in scenes requiring intimate object interactions (e.g., faceless cat in Fig. 4). In this work, we initialize holistic 3D representation through LLM-layout, which serves as a better warm-up for SDS optimization. The proposed TCSD further enhance optimization towards text-3D alignment. Our approach can generate holistic 3D content while facilitating more realistic interactions among various objects.
Multimodal Large Language Models.
Recent advancements in large language models (LLMs)[40, 29, 1] have led to increased interest in Multimodal Large Language Models (MLLMs), which combine vision understanding capabilities with traditional LLMs[29, 17]. Given their promising abilities in visual reasoning and understanding, MLLMs are being explored to enhance 3D generation, particularly in areas such as automatic evaluation [42, 10] and data preparation [38, 6]. However, existing approaches have not yet integrated MLLMs directly into the 3D generation process. In contrast, our work fine-tunes the LLaVA model [17] to provide semantic coherence criticism and view-checking based on multi-view images. We are the first to leverage MLLM guidance to assist directly in SDS optimization.
3 Preliminaries
Score Distillation Sampling.
Score Distillation Sampling (SDS) employs priors from pre-trained 2D diffusion models to facilitate the generation of 3D content, which is widely adopted in text-to-3D methods [30, 21, 4, 41, 25]. Given a user input prompt $y$, parameterized 3D representation $\theta$, and pre-trained 2D diffusion model $\Phi(\textbf{x}_{t}|y)$ along with noise prediction network $\epsilon_{\Phi}(\textbf{x}_{t},t,y)$, the objective of SDS is to optimize $\theta$ by minimizing the KL-divergence between the rendered image distribution $q_{t}^{\theta}(\textbf{x}_{t}|c)$ and sampled image distribution $p_{t}(\textbf{x}_{t}|y))$ from diffusion model:
$\mathcal{L}_{SDS}(\theta)=\min\limits_{\theta}D_{KL}(q_{t}^{\theta}(\textbf{x}_{t}|c)||p_{t}(\textbf{x}_{t}|y)).$ (1)
where $\textbf{x}_{t}$ represents the noisy rendered image $\textbf{x}=g(\theta,c)$ at timestamp $t$, and $g$ is the differentiable rendering function.
Ignoring the UNet Jacobian [30], the gradient computation of SDS loss is as follows:
$\small\begin{split}\nabla_{\theta}\mathcal{L}_{SDS}(\theta)&\triangleq\mathrm{E}_{t,\textbf{x}}[w(t)\frac{\sigma_{t}}{\alpha_{t}}\nabla_{\theta}D_{KL}(q_{t}^{\theta}(\textbf{x}_{t}|c,y)||p_{t}(\textbf{x}_{t}|y))]\\ &\triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\epsilon_{\Phi}(\textbf{x}_{t},t,y)-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\end{split}$ (2)
$\alpha_{t}$ and $\sigma_{t}$ are hyperparameters of noise schedule, $w(t)$ is the time-dependent weighting function, and $\epsilon_{\Phi}:=(1+s)\epsilon_{\Phi}(\textbf{x}_{t},t,y)-s\epsilon_{\Phi}(\textbf{x}_{t},t,\emptyset)$ is the modification of noise prediction with classifier-free guidance (CFG) as $s$.
4 Method
In this section, we introduce CoherenDream, a novel text-aligned 3D generation framework as depicted in Fig. 2. We first demonstrate the derivation of Textual Coherent Score Distillation (TCSD) in Sec. 4.1, which introduce MLLM feedback for across-view image distribution. Then we introduce a fine-tuned MLLM, 3DLLaVA-CRITIC to better evaluate across-view alignment in 3D generation in Sec. 4.2. Finally, we elaborate on the overall framework CoherenDream in Sec. 4.3, where we integrate three guidance tasks with TCSD and novel LLM-layout initialization technique to further enhance textual consistency.
4.1 Textual Coherent Score Distillation
In accordance with Eq. 1, we observe significant limitations in textual understanding and reasoning within the original SDS framework: (1) $p_{t}(\textbf{x}_{t}|y)$ is constrained by diffusion models, which may diverge considerably from the actual user-prompt distributions [16, 37]; (2) $p_{t}(\textbf{x}_{t}|y)$ heavily relies on the rendered image x of last updated 3D representation $\theta$, leading to bias accumulation. To address these limitations, we introduce informative feedback from MLLMs to $p_{t}(\textbf{x}_{t}|y)$ towards textual consistent distribution.
We first formulate the ideal textual consistent distribution $\hat{p}_{t}(\textbf{x}_{t}|y)$ as a joint distribution $\hat{p}_{t}(\textbf{x}_{t}|y)=p_{t}(\textbf{x}_{t}|y)p(y|\textbf{x}_{t})$. Based on the connection between diffusion models and score matching [36], the score function of $p_{t}(\textbf{x}_{t}|y)$ can be derived as:
$\nabla_{\textbf{x}_{t}}\text{log}\,p_{t}(\textbf{x}_{t}|y)=-\frac{1}{\sqrt{1-\alpha_{t}}}\epsilon_{\Phi}(\textbf{x},t,y)$ (3)
We substitute Eq. 3 into score function for $\hat{p}_{t}(\textbf{x}_{t}|y)$:
$\small\nabla_{\textbf{x}_{t}}\text{log}(p_{t}(\textbf{x}_{t}|y)p(y|\textbf{x}_{t}))=-\frac{1}{\sqrt{1-\alpha_{t}}}\epsilon_{\Phi}(\textbf{x},t,y)+\nabla_{\textbf{x}_{t}}\text{log}\,p(y|\textbf{x}_{t})$ (4)
Then, we derive the textual consistent noise prediction:
$\begin{split}\small\hat{\epsilon}_{\Phi}(\textbf{x}_{t},t,y)=\epsilon_{\Phi}(\textbf{x}_{t},t,y)-\sqrt{1-\alpha_{t}}\nabla_{\textbf{x}_{t}}\text{log}\,p(y|\textbf{x}_{t})\\ =\epsilon_{\Phi}(\textbf{x}_{t},t,y)+\lambda\cdot\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\textbf{x}_{t}))\end{split}$ (5)
where $\lambda=\sqrt{1-\alpha_{t}}$ controls the guidance strength, $\mathcal{L}_{ce}$ denotes autogressive cross-entropy loss, $f_{cr}$ indicates MLLM as guidance function, and T refers to predefined questions and answers related to $y$. However, deploying guidance on noisy input $\textbf{x}_{t}$ is not practical, we consider calculating feedback on one-step denoising data $\hat{\textbf{x}}_{0}$. Hence, the final feedback $\delta(\textbf{T},\hat{\textbf{x}}_{0})$ for textual consistency is represented as:
$\begin{split}\small\delta(\textbf{T},\hat{\textbf{x}}_{0})&=\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\hat{\textbf{x}}_{0}))\\ &=\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\frac{\textbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\Phi}(\textbf{x}_{t},t,y)}{\sqrt{\alpha}}))\end{split}\vspace{-3mm}$ (6)
Thus, the textual consistent gradient of Textual Coherent Score Distillation can be reformulated by Eq. 2 and ideal noise prediction in Eq. 5 as follows:
$\begin{split}\small\nabla_{\theta}\mathcal{L}_{TCSD}(\theta)\triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\hat{\epsilon}_{\Phi}(\textbf{x}_{t},t,y)-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\\ \triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\epsilon_{\Phi}(\textbf{x}_{t},t,y)+\lambda\cdot\delta(\textbf{T},\hat{\textbf{x}}_{0})-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\end{split}$ (7)
4.2 3DLLaVA-CRITIC
Original MLLMs are primarily designed as language assistants and cannot assess 3D generation adequately. To bridge this gap, we decompose the evaluation of textual consistency in 3D generation into three question-answering tasks: scene graph generation, multi-label image classification, and view classification. Building on this foundation, we introduce 3DLLaVA-CRITIC, following the architecture of LLaVA-OV [17] and fine-tuned using instruct-tuning pairs generated by advanced LLMs such as GPT-4o [29].
4.2.1 Task definition.
  • Scene graph generation. We establish this task to enhance global semantic understanding in across-view image distribution. Unlike caption generation, which often includes excessive and irrelevant descriptive details, the scene graph format requires the model to focus on object interactions. Specifically, given across-view observation, the 3DLLaVA-CRITIC produces a set of scene graphs represented as triplets. The form of scene graph triplet is [subject, relation, object] or [object, is, attribute]. To avoid ambiguity, we define four types of relation: Actions, Spatial Relations, Descriptive Verbs, and Non-specific Connections (e.g., “and”).
  • Multi-label classification. We observe that missing objects are a common phenomenon in textual inconsistency. Therefore, multi-label classification task asks 3DLLaVA-CRITIC to focus on primary objects presented in the user’s input. Specifically, 3DLLaVA-CRITIC need to extract objects in given observation.
  • View classification. Given input images, the 3DLLaVA-CRITIC determines the input camera position from the options: side, front, back, overhead. By comparing this classification with the sampled camera positions, the model guides the SDS to generate results that are consistent with the correct viewpoints.
Figure 3: View-aware data collection pipeline for 3DLLaVA-CRITIC that consist of (1) using LLM to generate diverse coarse text prompt and corresponding layout and rendering from random viewpoints in Blender; (2) random sampling images from T2I diffusion model conditioned on layout image and randomly assembling into a 2$\times$2 grid image; (3) employing GPT-4o to extract semantic annotation, including scene graph and view direction.
4.2.2 View-aware data collection pipeline.
Unlike natural images, the sampled image distribution in SDS is conditioned on noisy rendered images from randomly sampled camera poses. To achieve better assessment during SDS optimization, we introduce a view-aware data collection pipeline, as illustrated in Fig. 3. Detailed prompts for each step in the pipeline are available in the Appendix. Layout condition generation. Inspired by LayoutGPT [7], we utilize GPT-4 [29] to generate 200 diverse text prompts based on examples in the DreamFusion [30] library. Layout is defined as 3D bounding boxes with box center coordinates and dimension: (x, y, z, h, w, l). These layouts are rendered in  [^1] https://www.blender.org/Blender, and we randomly sample camera poses to produce 32 rendering images per layout.
View-aware image generation. Based on the previously created prompts and layout condition images, we enhance the original prompts by integrating view prompts (e.g., side, back, front, overhead). Utilizing DeepFloyd-IF [35] and MVDream [34], we generate view-aware images by varying random seeds, noise strengths, and denoising steps, in line with the optimization process of CoherenDream. Under each layout condition, we randomly select 4 camera views to create a 2$\times$2 grid image, resulting in a total of 36 grid images.
GPT-4o Annotation. After the last step, we curate 307,409 grid images and then prompt GPT-4o [29] to generate scene graphs, view prompts, and perform object extraction. These high-quality grid image-text pairs serve as instruction tuning data for proposed 3DLLaVA-CRITIC.
4.3 Framework of CoherenDream
Building upon the proposed TCSD optimization, we present the CoherenDream framework. This framework is based on NeRF [27] as 3D representation, utilizing Instant-NGP [28] with a volume renderer. To ensure coherence with user input, we incorporate three guidance tasks (as outlined in Sec. 4.2.1) with Eq. 7, which direct sampled image distribution toward the textual consistent distribution. During the optimization process, we employ well-established techniques such as time-annealing [14] and resolution scaling-up [41]. In addition, we introduce a novel technique called LLM-layout initialization, which significantly enhances the quality of the generated 3D content.
Guidance Tasks. We equipe TCSD optimization with three feedbacks from 3DLLaVA-CRITIC: in scene graph generation, multi-label classification and view classification. Detailed task descriptions can be found in Sec. 4.2.1. For the scene graph generation task, $\textbf{T}_{sg}$ can be extracted by humans or GPT according to user’s input. And answers $\textbf{T}_{ob}$ of multi-label classification can be directly obtained from $\textbf{T}_{sg}$. $\textbf{T}_{view}$ can be decided given the sampled camera pose. Therefore, the overall feedback of CoherenDream is demonstrated as: $\delta_{\text{CoherenDream}}=\lambda_{sg}\delta(\textbf{T}_{sg},\hat{\textbf{x}}_{0})+\lambda_{ob}\delta(\textbf{T}_{ob},\hat{\textbf{x}}_{0})+\lambda_{view}\delta(\textbf{T}_{view},\hat{\textbf{x}}_{0})$.
Figure 4: Qualitative comparison with representative methods. The results indicate that existing text-to-3D generation methods do not produce textual consistent results, involving objects omissions or unnatural interactions (highlighted in red). Conversely, our CoherenDream generates more textually faithful results that benefited from effective MLLM feedback.
LLM-layout Initialization. As discussed in Sec. 4.1, the rendered images x of 3D representations significantly influence the sampled image distribution from the diffusion model. The usage of a unit sphere for initialization provides less information condition, leading to the traption of local minimum during the score distillation optimization. It also leads to textual inconsistency. To address this issue, we introduce an LLM-generated layout to warm up CoherenDream generation, which we refer to as LLM-Layout. As illustrated in the “Layout Condition Generation” part of Fig. 3, LLM-Layout consists of a set of cubes in normalized space. The collection of LLM-Layouts mirrors the layout condition generation process detailed in Sec. 4.2.2. Specifically, we fit the density network of NeRF to align with the occupancy defined by the LLM-Layout. Since the LLM-Layout only provides fundamental placement information, we do not wish to constrain the 3D representation’s flexibility regarding shape. To balance this, we implement a decay in importance near the surface, allowing for more freedom in shape representation. The surface weight decay binary cross entropy loss function can be expressed as:
$\mathcal{L}_{\text{LLM-Layout}}=\mathcal{L}_{bce}(\text{occ}_{\theta}(p),\text{occ}_{\text{LLM-Layout}}(p))\dot{(}1-e^{\frac{d^{2}}{2\sigma}})$
where $\text{occ}_{\theta}$ indicates occupancy prediction from density network of NeRF $\theta$, $p$ is the points set sampled from camera view, $d$ denotes the distance of $p$ from the surface, and $\sigma$ is a hyperparameter that controls the strength of the constraint. During the warming-up phase, which consists of $N=600$ steps, the noising strength is restricted to the range $[0.4,0.7]$ to encourage the sampled distribution to adhere more closely to the initialized rendered image. Additionally, the importance of $\mathcal{L}_{\text{LLM-Layout}}$ decays over the warming-up steps. This approach enables us to achieve a better initialization for the SCSD optimization without severely restricting the representational capacity of NeRF.
[…]
MVDream [34] 39.0 28.5 57min
ProlificDreamer [41] 45.7 25.8 240min
RichDreamer [31] 34.8 22.0 70min
CoherenDream 39.4 29.1 60min
Table 1: Comparison on T3Bench multiple objects set.
Method CLIP Score $\uparrow$ TIFA Score $\uparrow$ VQAScore $\uparrow$
MVDream [34] 30.5 77.4 0.73
DreamView [44] 31.1 77.9 0.71
JointDreamer [15] 30.8 73.7 0.71
CoherenDream 31.4 81.4 0.79
Table 2: Comparison on TIFA v1.0 subset.
Figure 5: Comparing TCSD with Original SDS and GraphDreamer. Original SDS exhibits bias accumulation issues, resulting in object omissions. GraphDreamer, on the other hand, may encounter unnatural decomposition, leading to objects being fully obscured by others. In contrast, TCSD leverages a dynamic MLLM assessor to produce coherent results in a holistic 3D space.
5 Experiments
In this part, we present the text-to-3D generation results of CoherenDream with qualitative and quantitative evaluations, illustrating state-of-the-art performance. We also make an ablation analysis of the proposed TCSD. More details can be found in the Appendix.
5.1 Textual Consistent 3D Generation
Qualitative Comparison.
Figure 4 shows the comparison with several representative baselines, results produced by official codes. (i) MVDream [34]: While MVDream finetunes multi-view diffusion model on 3D dataset demonstrating heavy textual inconsistency with multi-object prompts, our CoherenDream triggers generalization of multi-view diffusion model guiding by MLLMs feedback. (ii) DreamView [44]: DreamView finetunes on 3D dataset leading to unrealistic texture and implausible interaction (e.g., the mixture of “bear” and “toy car”). Additionally, it need view-specific prompts, which cost human labor. (iii) JointDreamer [15]: JointDreamer maintains the generalization of original 2D diffusion, but it cannot break through the inherent limitations of diffusion model. It demonstrates object omissions (“child” in the third line) and insufficient semantic interaction understanding (targets for “beside” but shows “on” in the second row). In comparison, our Conherent can produce more text-faithful results. (iv) GALA3D [46]: GALA3D utilizes a compositional generative pipeline, optimizing each object separately. However, the non-overlapping layout leads to unnatural object placement, like a “party hat” above the “koala”. (v) GraphDreamer [9]: GraphDreamer advances compositional optimization through scene graph decomposition. However, it tends to have insufficient optimization and unnatural object reorganization. For example, faceless cat and incorrect understanding to “leans” in the last row. Note that the same layout and scene graph are used for GALA3D, GraphDreamer, and CoherenDream to ensure a fair comparison.
Figure 6: Incremental ablations on techniques in CoherenDream framework, which contributes to enhancing text alignment.
Figure 7: Ablation Study on Guidance Tasks with TCSD. Scene Graph Generation task provides valuable global semantic guidance, resulting in coherent geometric structures, although it may overlook some text-specific details. The Multi-label Classification task ensures correspondence between objects and text; however, it lacks the incorporation of attribution and relational information, leading to unrealistic structures. CoherenDream integrates these three guidance tasks to enhance both text alignment and geometric fidelity in 3D results.
Quantitative Comparison.
We perform quantitative evaluations on two benchmarks: the multi-object set from T3Bench [10] and a curated 45-prompt subset of TIFA v1.0 [13]. For T3Bench, we employ two complementary metrics: ImageReward [43] for assessing 3D asset quality and GPT-4 [42] for text-to-3D alignment verification. As shown in Table 2, our CoherenDream achieves comparable mesh and texture quality to the baseline MVDream [34], while demonstrating superior performance in alignment accuracy. For the TIFA subset, we adopt a multi-metric evaluation protocol to evaluate textual consistency: CLIP score [11], TIFA score [13] and VQAScore [23].To ensure comprehensive viewpoint coverage, we uniformly sample 10 azimuth angles around each 3D asset. For VQAScore, we report the maximum score across all viewpoints; For TIFA Score, we compute the intersection of correct answers across all viewpoints. For Clip Score, we adopt the CLIP ViT-B/32 as the feature extractor and calculate the average score across views. Quantitative results in Table 2 demonstrate the consistent superiority of CoherenDream in text-aligned 3D generation across all metrics. Specifically, CoherenDream achieves an improvement of the TIFA score by 4.0% over MVDream, demonstrating its superior corresponding to textual description benefiting from effective feedback from MLLM. Detailed examples of TIFA’s question-answering pairs can be found in the Appendix.
5.2 Ablation Study
More Comparison with GraphDreamer. As show in Fig. 5, we conduct further comparison with GraphDreamer [9]. For a fair comparison, our TCSD is only equiped with scene graph generation task and no LLM-layout initialization. The results demonstrate that GraphDreamer encounters unnatural reorganization, leading to objects being fully obscured by others. In contrast, our TCSD optimizes a holistic 3D representation and receives informative feedback from MLLMs.
Ablation on Textual Coherent Score Distillation. To verify the effectiveness of text-to-3D alignment achieved by SCSD, we distill from the multi-view diffusion model used in MVDream [34], which is known to overfit on 3D datasets. And we only utilize the original LLaVA-ov-0.5b with scene graph generation task as guidance. The primary distinction between the original SDS and SCSD, as illustrated in Fig. 1, is the introduction of an optimization direction $\delta$, which aims for ideal text alignment distribution defined in Eq. 6. As discussed in Sec. 4.1, the original SDS lacks the understanding and reasoning capabilities necessary for proper alignment, leading to significant misalignment issues. The incorporation of $\delta$ serves to mitigate these issues. The results presented in Fig. 5 validate our claims regarding the effectiveness of this approach.
Ablation on LLM-Layout Initialization and 3DLLaVA-CRITIC. We conduct incremental ablations on techniques in CoherenDream, focusing on LLM-Layout Initialization and the fine-tuned 3DLLaVA-CRITIC. As shown in Fig. 6, LLM-Layout Initialization enhances the warm-up process for the 3D representation, effectively activating semantic information within the diffusion model and preventing the optimization process from collapsing into local minima. Compared to the scenario without LLM-Layout, this technique helps the model focus on the primary objects and establishes reasonable relative scale relationships. Moreover, the proposed LLM-Layout Initialization strategy does not constrain the creativity of 3D presentations. However, due to the domain gap between natural images and the sampled image distribution from the diffusion model, the semantic feedback provided by the original MLLM does not accurately correspond to object attributes and secondary objects, such as ”a parrot is two” and ”tuxedo” in Fig. 6 are overlooked. In contrast, our fine-tuned 3DLLaVA-CRITIC offers improved text-to-3D understanding feedback, resulting in more text-faithful 3D content.
Ablation on Guidance Tasks. To discuss the effectiveness of the decomposed guidance tasks, we conduct studies based on the overall framework of CoherenDream, replacing the combined guidance tasks with individual ones. As shown in Fig. 7, the scene graph generation task provides global semantic feedback and ensures geometric fidelity. In contrast, the multi-label classification task focuses on the objects present in the image distribution but ignores their attributes and relationships, which can lead to unrealistic placements. The view classification task enhances 3D consistency but it is similar with multi-label classification, ignoring object’s attribute. Additionally, view classification requires encoding the camera pose as $\textbf{T}_{vc}$, resulting in slower training speeds. Our proposed CoherenDream integrates these three guidance tasks to enhance both text alignment and geometric fidelity in 3D contents.
6 Conclusion
In this paper, we introduce CoherenDream, a novel text-to-3D framework that leverages a powerful Multimodal Large Language Model (MLLM) to generate 3D results that are faithful to the user’s inputs. We demonstrate the framework’s effectiveness in optimizing score distillation through semantic feedback derived from the MLLM. Additionally, we incorporate LLM-Layout Initialization to enhance the warm-up process of the 3D representation. Our experiments show that CoherenDream achieves state-of-the-art performance in generating multi-object 3D content, excelling in both visual appearance and alignment with input text.